# Localize-and-Stitch: Efficient Model Merging via Sparse Task Arithmetic

### v1

@@ Abstract
Model merging offers an effective strategy to combine the strengths of multiple finetuned models into a unified model that preserves the specialized capabilities of each. Existing methods merge models in a global manner, performing arithmetic operations across all model parameters. However, such global merging often leads to task interference, degrading the performance of the merged model. In this work, we introduce Localize-and-Stitch, a novel approach that merges models in a localized way. Our algorithm works in two steps: i) Localization: identify tiny ($1\%$ of the total parameters) localized regions in the finetuned models containing essential skills for the downstream tasks, and ii) Stitching: reintegrate only these essential regions back into the pretrained model for task synergy. We demonstrate that our approach effectively locates sparse regions responsible for finetuned performance, and the localized regions could be treated as compact and interpretable representations of the finetuned models (tasks). Empirically, we evaluate our method on various vision and language benchmarks, showing that it outperforms existing model merging methods under different data availability scenarios. Beyond strong empirical performance, our algorithm also facilitates model compression and preserves pretrained knowledge, enabling flexible and continual skill composition from multiple finetuned models with minimal storage and computational overhead. Our code is available at https://github.com/yifei-he/Localize-and-Stitch.
@@ 1 Introduction
Pretrained models [14, 38, 50, 49] contain a wealth of rich and generalizable information, and finetuning these models for specific downstream tasks significantly enhances performance compared to training from scratch [5]. With the growing popularity of the pretrain-finetune paradigm, a vast array of finetuned models have been made available on platforms like Hugging Face [63], and many of them originate from the same pretrained models, such as CLIP [49]. However, deploying multiple finetuned models independently, each for a different downstream task, incurs large storage and maintenance cost, and limits knowledge transfer across them.
Model merging offers a viable solution to these challenges by integrating the strengths of multiple finetuned models into a single model that retains the specialized capabilities of each. The key advantage of model merging over traditional multi-task learning (MTL) [3, 72, 29, 24] is its efficiency, in that it does not require joint training on data across all tasks, but only involves arithmetic operations in the weight space. Existing methods merge models by averaging model parameters via arithmetic mean [64, 30], Fisher information [40], regression mean [32] or learned merging weights [68]. Those methods all average the models in a global manner, meaning that they perform arithmetic operations to all parameters of the finetuned models. However, similar to the conflicting gradient problem in MTL [71, 37], parameters in different finetuned models often have interference with each other, leading to suboptimal performance of the merged model. Recent works find that redundant parameter updates in finetuning are sources of conflicts [67]. Although the majority of model parameters are updated during finetuning, only very few contribute to improving the performance on downstream tasks [4, 27].
To overcome these limitations, we propose Localize-and-Stitch, an efficient algorithm that merges models in a localized manner. The algorithm (illustrated in Figure 1) involves two steps: i) Localization: identify tiny localized regions in the finetuned models containing essential skills for the downstream tasks. ii) Stitching: reintegrate only these essential regions back into the pretrained model. In the experiments, we verify that the changes in finetuend parameters are highly redundant, as we can efficiently identify just $1\%$ of the total parameters that recovers over $99\%$ of the finetuned performance. We evaluate our method on various language and vision tasks, showing that it outperforms existing model merging methods under different data availability scenarios.
@@ 2 Preliminaries
Task vectors. A task vector is the element-wise difference of the finetuned and pretrained parameters, denoted as $\tau_{i}=\theta_{\text{ft}}^{(i)}-\theta_{\text{pre}}\in\mathbb{R}^{d}$. These vectors encapsulate the knowledge acquired during the finetuning process. This knowledge can be effectively manipulated through task arithmetic [30], which involves performing arithmetic operations on task vectors to compose learned skills across tasks.
@@
Existing methods perform merging in the general form $\theta_{\text{merged}}=\theta_{\text{pre}}+\sum_{i=1}^{n}\alpha_{i}\tau_{i}$, and their difference mainly lies in the way of determining the scaling factors $\alpha_{i}$. We introduce the baselines in Table 1, and categorize them into global and localized methods based on whether the algorithm incorporates selection strategies to identify which parameters to merge. Localized algorithms specifically target sparse and localized regions, while global algorithms merge parameters indiscriminately. Note that AdaMerging has two variants: one learns layer-wise weights and another learns task-wise weights. In this work, we refer to AdaMerging as the layer-wise version because of its superior performance over its task-wise counterpart.
Data requirements. Fisher merging [40] requires over 256 data points per task to estimate the Fisher information. RegMean [32] requires more than 1600 data points per task to compute the inner product matrices effectively. AdaMerging [68] needs access to the full unlabeled test set for entropy minimization. In contrast, simple averaging [64], task arithmetic [30] and TIES-Merging [67] can be implemented without additional data. However, to achieve the best performance, both task arithmetic and TIES-Merging require tuning the hyperparameter $\alpha$ on a validation set.
@@
Global Simple averaging [64] $\theta_{\text{merged}}=\frac{1}{n}\sum_{i=1}^{n}\theta_{\text{ft}}^{(i)}$ Element-wise mean
Task arithmetic [30] $\theta_{\text{merged}}=\theta_{\text{pre}}+\alpha\sum_{i=1}^{n}\tau_{i}$ $\alpha$ tuned on a validation set
Fisher merging [40] $\theta_{\text{merged}}=\sum_{i=1}^{n}\hat{F}_{i}\theta_{\text{ft}}^{(i)}/\sum_{i=1}^{n}\hat{F}_{i}$ Weighted by Fisher information matrices
RegMean [32] $\theta_{\text{merged}}=(\sum_{i=1}^{n}X_{i}^{\top}X_{i})^{-1}\sum_{i=1}^{n}(X_{i}^{\top}X_{i}\theta_{\text{ft}}^{(i)})$ Minimizes difference in merged and individual activations
AdaMerging [68] $\{\theta_{\text{merged}}^{l}\}_{l=1}^{L}=\{\theta_{\text{pre}}+\sum_{i=1}^{n}\lambda_{i}^{l}\theta_{\text{ft}}^{(i)^{l}}\}_{l=1}^{L}$ Layer-wise weights learned by entropy minimization
Localized TIES-Merging [67] Trims the parameters in task vectors with small magnitudes, elect a sign at each position of the task vector and only keep the parameters with the same sign.
@@ 3.1 Motivation and objectives
Sparsity is important, but how to locate sparse regions is the key. Previous research identifies that during the finetuning stage, a significant portion of parameter updates is redundant, introducing interference in model merging [67]. This underscores the need for locating sparse regions to reduce such interference. While the importance of sparsity is recognized, strategies for achieving it remain underexplored. Earlier approaches typically identify sparse regions through random selection [70] or selecting regions with the top-$k\%$ largest magnitudes in task vectors [67]. However, they often fall short in identifying the most effective sparse regions for model merging. In Figure 3, we evaluate the efficacy of different localization methods across twelve language tasks, comparing the quality of their localized regions (specified by the binary mask $\gamma_{i}$). The performance is assessed on individual grafted models for each task, denoted as $\theta_{\text{pre}}+\gamma_{i}\odot\tau_{i}$, where $\tau_{i}$ is the task vector of the $i$-th task and $\odot$ is the element-wise product. This grafted performance measures how well the finetuned skills are preserved when only keeping parameter updates during finetuning in localized regions. Unlike previous methods, we directly optimize the binary masks to maximally retain finetuned performance, detailed in Section 3.2. Our method significantly outperforms others, especially at lower sparsity levels. The strength of our approach lies in its precision in identifying small but informative regions, which is particularly advantageous for model merging.
@@
Sparse regions with less overlap reduce task conflicts. Identifying the smallest possible regions with essential finetuned skills is key to minimizing potential conflicts among task vectors, as smaller localized regions naturally incur less overlap among tasks. With reduced overlap, each task can occupy its own, relatively disjoint localized region, thereby reducing task conflicts. This has the intuitive explanation that when two conflicting tasks share highly overlapping localized regions, they will compete to steer the parameters within these regions to their advantage in the merged model, leading to performance degradation. We demonstrate this by a case study (Figure 3) on merging models finetuned on two conflicting language tasks: QNLI [59] and MNLI [2]. QNLI involves predicting whether a context contains the answer to the given question, and MNLI involves predicting text entailment given a sentence pair. These tasks are conflicting, manifested by a noticeable performance decline for both tasks when using simple averaging to merge the corresponding finetuned models. However, if the localized regions are small yet sufficiently informative about their respective tasks, the reduced overlap between these regions decreases task conflicts and enhances overall performance after merging. In other words, as long as the localized region contains sufficient task-specific knowledge, including more parameters than necessary in them only introduces additional task interference.

### v2

# algorithmic

@@ Abstract
Model merging offers an effective strategy to combine the strengths of multiple finetuned models into a unified model that preserves the specialized capabilities of each. Existing methods merge models in a global manner, performing arithmetic operations across all model parameters. However, such global merging often leads to task interference, degrading the performance of the merged model. In this work, we introduce Localize-and-Stitch, a novel approach that merges models in a localized way. Our algorithm works in two steps: i) Localization: identify tiny ($1\%$ of the total parameters) localized regions in the finetuned models containing essential skills for the downstream tasks, and ii) Stitching: reintegrate only these essential regions back into the pretrained model for task synergy. We demonstrate that our approach effectively locates sparse regions responsible for finetuned performance, and the localized regions could be treated as compact and interpretable representations of the finetuned models (tasks). Empirically, we evaluate our method on various vision and language benchmarks, showing that it outperforms existing model merging methods under different data availability scenarios. Beyond strong empirical performance, our algorithm also facilitates model compression and preserves pretrained knowledge, enabling flexible and continual skill composition from multiple finetuned models with minimal storage and computational overhead. Our code is available at https://github.com/uiuctml/Localize-and-Stitch.
@@ 1 Introduction
Pretrained models (Devlin et al., 2018; Liu et al., 2019; Raffel et al., 2020; Radford et al., 2021) contain a wealth of rich and generalizable information, and finetuning these models for specific downstream tasks significantly enhances performance compared to training from scratch (Chen et al., 2020b). With the growing popularity of the pretrain-finetune paradigm, a vast array of finetuned models have been made available on platforms like Hugging Face (Wolf et al., 2020), and many of them originate from the same pretrained models, such as CLIP (Radford et al., 2021). However, deploying multiple finetuned models independently, each for a different downstream task, incurs large storage and maintenance cost, and limits knowledge transfer across them.
Model merging offers a viable solution to these challenges by integrating the strengths of multiple finetuned models into a single model that retains the specialized capabilities of each. The key advantage of model merging over traditional multi-task learning (MTL) (Caruana, 1997; Zhang & Yang, 2021; Hu et al., 2024; He et al., 2024) is its efficiency, in that it does not require joint training on data across all tasks, but only involves arithmetic operations in the weight space. Existing methods merge models by averaging model parameters via arithmetic mean (Wortsman et al., 2022a; Ilharco et al., 2023), Fisher information (Matena & Raffel, 2022), regression mean (Jin et al., 2022) or learned merging weights (Yang et al., 2023). Those methods all average the models in a global manner, meaning that they perform arithmetic operations to all parameters of the finetuned models. However, similar to the conflicting gradient problem in MTL (Yu et al., 2020; Liu et al., 2021), parameters in different finetuned models often have interference with each other, leading to suboptimal performance of the merged model. Recent works find that redundant parameter updates in finetuning are sources of conflicts (Yadav et al., 2023). Although the majority of model parameters are updated during finetuning, only very few contribute to improving the performance on downstream tasks (Chen et al., 2020a; Hoefler et al., 2021).
To overcome these limitations, we propose Localize-and-Stitch, an efficient algorithm that merges models in a localized manner. The algorithm (Figure 1) involves two steps: i) Localization: identify tiny localized regions in the finetuned models containing essential skills for the downstream tasks. ii) Stitching: reintegrate only these essential regions back into the pretrained model. In the experiments, we verify that the changes in finetuend parameters are highly redundant, as we can efficiently identify just $1\%$ of the total parameters that recovers over $99\%$ of the finetuned performance. We evaluate our method on various language and vision tasks, showing that it outperforms existing model merging methods under different data availability scenarios.
@@ 2 Preliminaries
Task vectors. A task vector is the element-wise difference of the finetuned and pretrained parameters, denoted as $\tau_{i}=\theta_{\text{ft}}^{(i)}-\theta_{\text{pre}}\in\mathbb{R}^{d}$. These vectors encapsulate the knowledge acquired during the finetuning process. This knowledge can be effectively manipulated through task arithmetic (Ilharco et al., 2023), which involves performing arithmetic operations on task vectors to compose learned skills across tasks.
@@
Existing methods perform merging in the general form $\theta_{\text{merged}}=\theta_{\text{pre}}+\sum_{i=1}^{n}\alpha_{i}\tau_{i}$, and their difference mainly lies in the way of determining the scaling factors $\alpha_{i}$. We introduce the baselines in Table 1, and categorize them into global and localized methods based on whether the algorithm incorporates selection strategies to identify which parameters to merge. Localized algorithms specifically target sparse and localized regions, while global algorithms merge parameters indiscriminately. We provide detailed comparisons with the two localized algorithms, in Appendix C and Appendix D respectively. Note that AdaMerging has two variants: one learns layer-wise weights and another learns task-wise weights. In this work, we refer to AdaMerging as the layer-wise version because of its superior performance over its task-wise counterpart.
Data requirements. Fisher merging (Matena & Raffel, 2022) requires over 256 data points per task to estimate Fisher information. RegMean (Jin et al., 2022) requires more than 1600 data points per task to compute the inner product matrices effectively. AdaMerging (Yang et al., 2023) needs access to the full unlabeled test set for entropy minimization. Consensus TA/TIES (Wang et al., 2024b) requires a validation set to tune hyperparameters. In contrast, simple averaging (Wortsman et al., 2022a), task arithmetic (Ilharco et al., 2023) and TIES-Merging (Yadav et al., 2023) can be implemented without additional data. However, to achieve the best performance, both task arithmetic and TIES-Merging require tuning the hyperparameter $\alpha$ on a validation set.
@@
Global Simple averaging (Wortsman et al., 2022a) $\theta_{\text{merged}}=\frac{1}{n}\sum_{i=1}^{n}\theta_{\text{ft}}^{(i)}$ Element-wise mean
Task arithmetic (TA) (Ilharco et al., 2023) $\theta_{\text{merged}}=\theta_{\text{pre}}+\alpha\sum_{i=1}^{n}\tau_{i}$ $\alpha$ tuned on a validation set
Fisher merging (Matena & Raffel, 2022) $\theta_{\text{merged}}=\sum_{i=1}^{n}\hat{F}_{i}\theta_{\text{ft}}^{(i)}/\sum_{i=1}^{n}\hat{F}_{i}$ Weighted by Fisher information matrices
RegMean (Jin et al., 2022) $\theta_{\text{merged}}=(\sum_{i=1}^{n}X_{i}^{\top}X_{i})^{-1}\sum_{i=1}^{n}(X_{i}^{\top}X_{i}\theta_{\text{ft}}^{(i)})$ Minimizes difference in merged and individual activations
AdaMerging (Yang et al., 2023) $\{\theta_{\text{merged}}^{l}\}_{l=1}^{L}=\{\theta_{\text{pre}}+\sum_{i=1}^{n}\lambda_{i}^{l}\theta_{\text{ft}}^{(i)^{l}}\}_{l=1}^{L}$ Layer-wise weights learned by entropy minimization
Localized TIES-Merging (Yadav et al., 2023) Trims the parameters in task vectors with small magnitudes, elect a sign at each position of the task vector and only keep the parameters with the same sign.
Consensus TA/TIES (Wang et al., 2024b) Compute multi-task task vectors: $\tau_{MTL}=\theta_{\text{merged}}-\theta_{\text{pre}}$ with $\theta_{\text{merged}}$ obtained by any merging method. Construct task masks: $m_{t}=\mathbbm{1}\{|\tau_{t}|\geq|\tau_{MTL}-\tau_{t}|\cdot\lambda_{t}\}$. Apply consensus mask $m_{consensus}=\mathbbm{1}\{\sum_{t\in[T]}m_{t}\geq 2\}$ on $\tau_{MTL}$.
@@ 3.1 Motivation and objectives
Sparsity is important, but how to locate sparse regions is the key. Previous research identifies that during the finetuning stage, a significant portion of parameter updates is redundant, introducing interference in model merging (Yadav et al., 2023). This underscores the need for locating sparse regions to reduce such interference. While the importance of sparsity is recognized, strategies for achieving it remain underexplored. Earlier approaches typically identify sparse regions through random selection (Yu et al., 2023) or selecting regions with the top-$k\%$ largest magnitudes in task vectors (Yadav et al., 2023). However, they often fall short in identifying the most effective sparse regions for model merging. In Figure 3, we evaluate the efficacy of different localization methods across twelve language tasks, comparing the quality of their localized regions (specified by the binary mask $\gamma_{i}$). The performance is assessed on individual grafted models for each task, denoted as $\theta_{\text{pre}}+\gamma_{i}\odot\tau_{i}$, where $\tau_{i}$ is the task vector of the $i$-th task and $\odot$ is the element-wise product. This grafted performance measures how well the finetuned skills are preserved when only keeping parameter updates during finetuning in localized regions. Unlike previous methods, we directly optimize the binary masks to maximally retain finetuned performance, detailed in Section 3.2. Our method significantly outperforms others, especially at lower sparsity levels. The strength of our approach lies in its precision in identifying small but informative regions, which is particularly advantageous for model merging.
@@
Sparse regions with less overlap reduce task conflicts. Identifying the smallest possible regions with essential finetuned skills is key to minimizing potential conflicts among task vectors, as smaller localized regions naturally incur less overlap among tasks. With reduced overlap, each task can occupy its own, relatively disjoint localized region, thereby reducing task conflicts. This has the intuitive explanation that when two conflicting tasks share highly overlapping localized regions, they will compete to steer the parameters within these regions to their advantage in the merged model, leading to performance degradation. We demonstrate this by a case study (Figure 3) on merging models finetuned on two conflicting language tasks: QNLI (Wang et al., 2018) and MNLI (Bowman et al., 2015). QNLI involves predicting whether a context contains the answer to the given question, and MNLI involves predicting text entailment given a sentence pair. These tasks are conflicting, manifested by a noticeable performance decline for both tasks when using simple averaging to merge the corresponding finetuned models. However, if the localized regions are small yet sufficiently informative about their respective tasks, the reduced overlap between these regions decreases task conflicts and enhances overall performance after merging. In other words, as long as the localized region contains sufficient task-specific knowledge, including more parameters than necessary in them only introduces additional task interference.
@@ 3.2 Localization
The objectives are grounded in the findings of Panigrahi et al. (2023), which demonstrates that skills developed during finetuning are localizable. Specifically, grafting a small subset of finetuned parameters onto the pretrained model can almost fully recover the performance of the finetuned model. Panigrahi et al. (2023) propose the following optimization problem to identify the localized parameters, and we adapt it in the model merging setting. With the constraint on the sparsity level $s$, on the $i$-th task with the loss function $\ell_{i}$ and task vector $\tau_{i}$, we optimize for a binary mask $\gamma_{i}$ such that only adding up the masked portion of the task vector onto the pretrained model performs well on the $i$-th task
$\gamma_{i}=\operatorname*{arg\,min}_{\gamma\in\{0,1\}^{d}:\|\gamma\|_{0}\leq s}\ell_{i}(\theta_{\text{pre}}+\gamma\odot\tau_{i}),$
where $\odot$ denotes the element-wise product. For the ease of optimization, we follow Panigrahi et al. (2023) to reparametrize the binary mask $\gamma$ as a real-valued vector $S$. To control the sparsity, we additionally relax the $L_{0}$ sparsity constraint to be $L_{1}$. As a result, the optimization is reformulated as
$S_{i}=\operatorname*{arg\,min}_{S\in\mathbb{R}^{d}}\ell_{i}\left(\theta_{\text{pre}}+\sigma(S)\odot\tau_{i}\right)+\lambda\|\sigma(S)\|_{1},$ (1)
where $\sigma$ is the sigmoid function, and $\lambda$ controls the strength of the $L_{1}$ regularization. At the end of the optimization, we round $\sigma(S_{i})$ to be binary.
In comparison, Panigrahi et al. (2023) uses the following formulation
$S_{i}=\operatorname*{arg\,min}_{S\in\mathbb{R}^{d}}\ell_{i}(\gamma\odot\theta_{\text{ft}}+(1-\gamma)\odot\theta_{\text{pre}}),$
where $\gamma_{base}$ is the top-$k\%$ largest elements in the task vector, which serves as an initialization for the optimization. There are two main advantages of formulation 1 over 3.2. Firstly, our formulation of $S$ is more straightforward, as we directly have $\gamma=\sigma(S)$. In contrast, $S$ in Section 3.2 serves as a selector of whether to take the value from $\gamma_{base}$, leading to more complex computation. Secondly, our approach uses the $L_{1}$ constraint to control the sparsity in a more fine-grained manner, while Section 3.2 does not have this constraint, and they control the sparsity via early stopping instead. A detailed empirical performance comparison between our localization technique and the one in (Panigrahi et al., 2023) is presented in Appendix E.
Note that the optimization is highly efficient, requiring as few as 8-shot data with 10 epochs of training using SGD. A detailed ablation on the impact of data availability on the mask quality is shown in Section 4.4.
(a) Jaccard similarity of pairwise task masks.
(b) Cosine similarity of masked task vectors.
[…]
Interpretation of task relationships.
In Figure 4, we validate that our localization method effectively identifies task-specific regions with minimal overlap. In Figure 4(a), we report the Jaccard similarity of all pairwise task masks, namely for each pair of masks $\gamma_{i}$ and $\gamma_{j}$, we compute $|\gamma_{i}\cap\gamma_{j}|/|\gamma_{i}\cup\gamma_{j}|$. The majority of task pairs exhibit a Jaccard similarity below $5\%$, confirming minimal overlap. For the few pairs with Jaccard similarity larger than $10\%$ (upper left corner of Figure 4(a)), we further compute the cosine similarity of their masked task vectors in Figure 4(b), and find that their cosine similarities are almost 1, indicating high agreement of the parameters within the overlapped regions. Since these four tasks are all sentiment classification tasks, this phenomenon intuitively suggests a shared skill set across the tasks, located in the overlapped regions. We elaborate our resolution for the overlapped regions in Section 3.3
Distribution of localized regions.
We analyze the distribution of the localized regions for language tasks in RoBERTa-base models Figure 5, both in terms of the layer index and the transformer components. For the layers, different tasks seem to occupy different layers, although the earlier layers in the network seldomly appear in the localized regions. Interestingly, most of the localized regions concentrate in the LayerNorm (Ba, 2016) parameters. This pattern can possibly be attributed to a distribution shift observed in the finetuning data compared to the pretraining data, necessitating adjustments to the LayerNorm parameters to accommodate this shift. The same plots for GPT2-XL and ViT can be found in Appendix B, and the findings hold true for those models as well.
Localization without validation data.
In the rare cases where no labeled data is available, we adopt a similar strategy as the “Trim” step in TIES-Merging (Yadav et al., 2023), which selects positions in task vectors with the top-$k\%$ largest magnitudes, i.e., the parameters changed the most during finetuning. We refer to our approach as Dataless Localize-and-Stitch. As shown in Figure 3, to match the performance of localization with validation data, the dataless version typically requires locating larger regions. This expansion is necessary to encapsulate sufficient skills acquired during finetuning, but it also leads to increased task conflicts. Nevertheless, in Section 4, we show that our dataless version still outperforms all other methods that do not require additional validation data.
Despite the similarity, there are two key differences between Dataless Localize-and-Stitch and TIES: i) Smaller localized region: We find selecting the top-$5\%$ of parameters is sufficient for our pipeline, compared to the top-$20\%$ recommended by TIES-Merging. Our smaller selected region incurs less overlapping, leading to reduced task interference. Note that this is not the only advantage of our approach, as reducing the threshold in TIES to be $5\%$ does not yield an improved performance as demonstrated in Appendix C (Tables 12 and 13). ii) Better merging performance: We use “Stitching” described in the next section for merging the localized regions, instead of the “Elect” procedure in TIES. The “Elect” approach in TIES might work well when overlapping regions involve a larger number of tasks, allowing sign determination and selective averaging to better capture a consensus among task vectors. However, when only two tasks are involved in the overlapping regions (which is often the case as shown in Figure 13 in Appendix C), TIES may only retain parameters predominantly from the task with the larger magnitude at each position. In such scenarios, important parameters for both tasks could be alternately ignored, impairing the overall efficacy in maintaining crucial task-specific information, particularly in tightly contested regions. We provide a more detailed discussion about the advantages of Dataless Localize-and-Stitch over TIES-Merging with empirical evidence in Appendix C.
3.3 Stitching
After obtaining the binary mask for each task, we integrate these masks, and apply them to task vectors to construct the merged model. Given the sparsity, the masks generally activate different positions for different tasks, minimizing overlap. However, in instances where overlaps occur – that is, where multiple tasks share the same mask positions – we address this by averaging the parameters in these regions. Specifically, for each final mask $\gamma_{i}^{\prime}$, the value at the $k$-th position, denoted $\gamma_{i}^{\prime}[k]$, is calculated as the reciprocal of the total number of tasks that have a mask value of 1 at that position; if the original mask value $\gamma_{i}[k]$ is 0, it remains 0 in the processed mask: $\gamma_{i}^{\prime}[k]=\gamma_{i}[k]\bigg{/}\left(\sum_{j=1}^{n}\gamma_{j}[k]\right).$ After we obtain the processed masks $\{\gamma_{i}^{\prime}\}_{i=1}^{n}$, we apply them to the task vectors and stitch the masked task vectors to get the final merged model